# Better bounds on Grothendieck constants of finite orders

### v2

@@ Abstract
Grothendieck constants $K_{G}\left({d}{}\right)$ bound the advantage of $d$-dimensional strategies over $1$-dimensional ones in a specific optimisation task. They have applications ranging from approximation algorithms to quantum nonlocality. However, apart from $d=2$, their values are unknown. Here, we exploit a recent Frank-Wolfe approach to provide good candidates for lower bounding some of these constants. The complete proof relies on solving difficult binary quadratic optimisation problems. For $d\in\{3,4,5\}$, we construct specific rectangular instances that we can solve to certify better bounds than those previously known; by monotonicity, our lower bounds improve on the state of the art for $d\leqslant 9$. For $d\in\{4,7,8\}$, we exploit elegant structures to build highly symmetric instances achieving even greater bounds; however, we can only solve them heuristically. We also recall the standard relation with violations of Bell inequalities and elaborate on it to interpret generalised Grothendieck constants $K_{G}\left({d}\mapsto{2}\right)$ as the advantage of complex quantum mechanics over real quantum mechanics. Motivated by this connection, we also improve the bounds on $K_{G}\left({d}\mapsto{2}\right)$.
@@ I Introduction
Quantum information is another field where this result is influential: following early observations by Tsirelson Tsi87 , an explicit connection has been established with the noise robustness in Bell experiments Bel64 ; AGT06 . These experiments aim at exhibiting a fascinating property of quantum mechanics in correlation scenarios: nonlocality BCP+14 . The link with Grothendieck’s theorem has then raised a surge of interest for the value of $K_{G}\left({3}{}\right)$, the Grothendieck constant of order three. Many works have thus demonstrated increasingly precise lower bounds CHSH69 ; Ver08 ; HLZ+15 ; BNV16 ; DBV17 ; DIB+23 and upper bounds Kri79 ; HQV+17 ; DIB+23 on its value. More recently, a numerical method has also been developed to come up with an even more precise (but not provable) estimate of $K_{G}\left({3}{}\right)$ SM24 .
@@
In this article, we combine the recent projection technique from DIB+23 with the powerful solver developed in DBV17 to obtain better lower bounds on $K_{G}\left({d}{}\right)$ for $3\leqslant d\leqslant 9$. Following FR94 , we also consider symmetric structures in high dimensions emerging from highly symmetric line packings FJM18 to suggest even better bounds on $K_{G}\left({4}{}\right)$, $K_{G}\left({7}{}\right)$, and $K_{G}\left({8}{}\right)$, that we unfortunately cannot prove as they involve optimisation problems that we only solve heuristically. We also consider the generalised Grothendieck constants $K_{G}\left({d}\mapsto{2}\right)$ and interpret them as the advantage of $d$-dimensional quantum mechanics over real quantum mechanics, a fact that was already studied in VP08 ; VP09 , but recently received more attention through the powerful results of RTW+21 . Our bounds on $K_{G}\left({d}{}\right)$ are analytical, while those on $K_{G}\left({d}\mapsto{2}\right)$ strongly rely on numerical methods: we post-process the upper bound on $K_{G}\left({3}\mapsto{2}\right)$ to convert it into an exact result, but the lower bounds remain inaccurate.
@@ II Preliminaries
$\mathrm{SDP}_{d}\left({M}\right)=\max\left\{\sum_{x=1}^{m_{1}}\sum_{y=1}^{m_{2}}M_{xy}\left\langle{a_{x}},{b_{y}}\right\rangle\leavevmode\nobreak\ \Big{|}\leavevmode\nobreak\ \forall x\in[m_{1}],\leavevmode\nobreak\ a_{x}\in S^{d-1},\leavevmode\nobreak\ \forall y\in[m_{2}],\leavevmode\nobreak\ b_{y}\in S^{d-1}\right\},$ (1)
where $[m]=\{1,\ldots,m\}$ and where $S^{d-1}$ is the $d$-dimensional unit sphere, which reduces to $S^{0}=\{-1,1\}$ for $d=1$. By introducing the set
$\mathrm{SDP}_{d}^{m_{1},m_{2}}=\operatorname{conv}\left\{X\leavevmode\nobreak\ \big{|}\leavevmode\nobreak\ X_{xy}=\left\langle{a_{x}},{b_{y}}\right\rangle,\leavevmode\nobreak\ \forall x\in[m_{1}],\leavevmode\nobreak\ a_{x}\in S^{d-1},\leavevmode\nobreak\ \forall y\in[m_{2}],\leavevmode\nobreak\ b_{y}\in S^{d-1}\right\},$ (2)
@@
$\mathrm{SDP}_{d}\left({M}\right)=\max\left\{\left\langle{M},{X}\right\rangle\leavevmode\nobreak\ \big{|}\leavevmode\nobreak\ X\in\mathrm{SDP}_{d}^{m_{1},m_{2}}\right\}.$ (3)
@@
$K_{G}\left({d}\mapsto{n}\right)=\sup\left\{\frac{\mathrm{SDP}_{d}\left({M}\right)}{\mathrm{SDP}_{n}\left({M}\right)}\leavevmode\nobreak\ \Big{|}\leavevmode\nobreak\ M\in\mathbb{R}^{m_{1}\times m_{2}},\quad m_{1},m_{2}\in\mathbb{N}\right\}.$ (5)
The standard Grothendieck constant of order $d$ is obtained when $n=1$, in which case we use the shorthand notation $K_{G}\left({d}\mapsto{1}\right)=K_{G}\left({d}{}\right)$. The Grothendieck constant of infinite order $K_{G}$ initially studied in Gro53 corresponds to the limit $\lim_{d\rightarrow\infty}K_{G}\left({d}{}\right)$, but it is out of the scope of our work, although we briefly mention it in Section V. We refer the reader interested in general aspects and further generalisations of these constants to Bri11 .
@@ III.1 Obtaining facets of the symmetrised correlation polytope
When $n=1$, the set $\mathrm{SDP}_{1}$ is a polytope, called the correlation polytope Pit91 . In BNV16 ; DBV17 ; DIB+23 the method used to solve 1 is to start from a point $P\in\mathrm{SDP}_{d}$ and to derive $M$ as a hyperplane separating $P$ from $\mathrm{SDP}_{1}$. This hyperplane is obtained by solving the projection of $P$ onto the correlation polytope $\mathrm{SDP}_{1}$ via Frank-Wolfe algorithms. We refer to Pok23 for a gentle introduction to FW algorithms, to BCC+22 for a complete review, and to (DIB+23, , Appendix C) for the details of our implementation. Note that some accelerations developed in DIB+23 ; DVP24 and used in this work are now part of the FrankWolfe.jl package BDH+24 .
In particular, the symmetrisation described in DVP24 is crucial, but as it depends on the underlying group and its action of the matrices we consider, exposing the full structure of the symmetrised correlation polytope for each group would be tedious. Instead, we give an brief general definition of this polytope and refer to DVP24 for a detailed example. Given a group $G$ acting on $[m_{1}]$ and $[m_{2}]$ (by means of signed permutations), the symmetrised correlation polytope $G(\mathrm{SDP}_{1}^{m_{1},m_{2}})$ is the convex hull of the averages of all orbits of the vertices of $\mathrm{SDP}_{1}^{m_{1},m_{2}}$ under the action of $G$. Note that this polytope lives in a subspace of the space invariant under the action of $G$ on matrices of size $m_{1}\times m_{2}$.
In Fig. 1 we illustrate the various geometrical cases that we encounter in this work. Their understanding justifies the procedure that we describe in Algorithm 1 to derive facets of the symmetrised correlation polytope, that is, separating hyperplanes touching the polytope on a space of codimension one. Importantly, for symmetric instances, the dimension of the ambient space is strictly smaller than $m_{1}m_{2}$. In the following, we consistently denote these facets by $A$, while $M$ will indicate separating hyperplanes without this extra property. Putative facets will also be denoted by $A$, that is, in cases where we rely on heuristic methods to compute $\mathrm{SDP}_{1}\left({A}\right)$.
These heuristic methods play an important role throughout the FW algorithm as they quickly give a good direction to make primal progress. Here we present them in a generalised framework that will turn useful when generalising our algorithm to $K_{G}\left({d}\mapsto{2}\right)$ in Section IV. At a given step of the FW algorithm minimising the squared distance to the set $\mathrm{SDP}_{n}$, finding the best direction with respect to the current gradient $M$ is exactly the problem in Eq. 3. This subroutine is called the Linear Minimisation Oracle (LMO) and, in the course of the algorithm, we usually use an alternating minimisation to obtain heuristic solutions that are enough to make progress (DIB+23, , Appendix B.1).
@@
Figure 1: Illustration of our projection method to separate a matrix $P\in\mathrm{SDP}_{d}$ from the correlation polytope $\mathrm{SDP}_{1}$. (a) In all cases that we consider in this work, the direction $P$ already defines, together with the corresponding value of $\mathrm{SDP}_{1}\left({P}\right)$, a separating hyperplane. (b) When running our FW algorithm to minimise the distance between $P$ and $\mathrm{SDP}_{1}$, the orthogonal projection obtained does not, in general, lie on a facet. The resulting separating hyperplane $M$ is better than $P$ from (a), but still not optimal. Note that this situation can also happen when starting from a point $vP$ on the line $[0,P]$ that is not close enough to $\mathrm{SDP}_{1}$ (not depicted here). (c) Even when $v$ is chosen small enough, if our FW algorithm has not fully converged to the orthogonal projection of $vP$, the separating hyperplane $M$ slightly deviates from a facet. (d) For a small enough $v$ and when our FW algorithm has converged, the resulting separating hyperplane is a facet, denoted $A$, that is, the tightest separation leading to the best value of $\left\langle{A},{P}\right\rangle/\mathrm{SDP}_{1}\left({A}\right)$.
@@
2:$v\leftarrow 1$triangleright $P$ assumed outside of $\mathrm{SDP}_{1}^{m_{1},m_{2}}$, see Fig. 1a
3:$\mathcal{S}\leftarrow\varnothing$triangleright empty active set to start with
@@
5:     $X,\mathcal{S}\leftarrow\mathrm{BPCG}(vP)$triangleright $X$ is the projection of $vP$ onto $\mathrm{SDP}_{1}^{m_{1},m_{2}}$ obtained with Algorithm 2 from DIB+23
6:     $M\leftarrow vP-X$triangleright direction of the separating hyperplane given by the final gradient
7:     $v\leftarrow\left\langle{M},{P}\right\rangle/\mathrm{SDP}_{1}\left({M}\right)$triangleright new point on the line $[0,P]$, strictly closer to $\mathrm{SDP}_{1}^{m_{1},m_{2}}$ than the previous one
8:end $\triangleright$ $\operatorname{codim}\mathcal{S}=1$, the ambient dimension may be smaller than $m_{1}m_{2}$ by symmetry
9:Output: $A\perp\mathcal{S}$triangleright normal vector defining the facet $A$ of the symmetrisation of $\mathrm{SDP}_{1}^{m_{1},m2}$ BBB+12 ; DVP24
@@ III.2 Heuristic results with highly symmetric line packings
Among the structures that we present in Table 1, the $d$-dimensional configurations reaching the best known kissing number in dimension $d$ are of particular interest. In particular, in dimension $d=3$, different realisations of the kissing number give rise to different values of $\mathrm{SDP}_{3}\left({A}\right)/\mathrm{SDP}_{1}\left({A}\right)$. As the icosahedron is an optimal line packing, this indicates that giving good bounds on $K_{G}\left({d}{}\right)$ does not, in general, boil down to finding good line packings. Note that the unicity of the configuration in dimension $d=4$ was recently proven LLM24 .
Upon inspection of Table 1 it is apparent that, in the small dimensions that we consider in this work ($d\leqslant 8$), root systems play an important role in obtaining good kissing configurations. These structures are inherently symmetric and play an important role in the theory of Lie groups and Lie algebras. We refer to Bak02 for definitions and elementary properties. The family $D_{d}$ has been studied in FR94 ; there, starting from the Gram matrix $P$ of size $d(d-1)$, the exact value of $\mathrm{SDP}_{1}\left({P}\right)$ is computed by invoking symmetry arguments. Moreover, for $d\in\{3,4,5\}$, the optimal diagonal modification of $\lambda=\frac{2}{3}$ is derived for $D_{d}$; in other words, FR94 shows that, among all matrices with shape $P-\lambda\mathds{1}$, the matrix $P-\frac{2}{3}\mathds{1}$ is giving the best lower bound on $K_{G}\left({d}{}\right)$. Here we go a bit further by extending this result to $d\in\{6,7,8\}$, a fact that was suspected in FR94 , but we also observe that the corresponding matrices $A=P-\frac{2}{3}\mathds{1}$ are actually facets of the symmetrised polytope $D_{d}(\mathrm{SDP}_{1}^{d(d-1)})$.
Moreover, we computationnally establish optimal diagonal modifications for other configurations, see Table 1. About the root system $E_{8}$, we note that the value $45/31$ reached without diagonal modification is already provided in (FR94, , page 50), where it is attributed to Reeds and Sloane. It indeed follows from (FR94, , Lemma 2) and from the transitivity of the Weyl group, see the acknowledgements in Section VII. However, these arguments do not directly apply to solve the diagonally modified case.
Note that in all cases involving irrational numbers (except the icosahedron), the facet $A$ cannot be obtained via a diagonal modification of $P$. This is because the off-diagonal elements of $P$ feature rational and irrational numbers. The exception of the icosahedron is due to its extra property of being an equiangular tight frame (ETF), all off-diagonal elements being ±$\pm 1/\sqrt{5}$, which can then lead to a facet by taking an irrational diagonal modification.
@@
[…]
$\frac{9-3\sqrt{5}}{2}$ $\approx 1.1459\hphantom{{}^{*}}$ PV22 $\frac{1+3\sqrt{5}}{6}$ $\approx 1.2847$ Icosahedron x $H_{3}$  $\frac{15-\sqrt{5}}{15}$
@@
5 20 $\frac{4}{3}$ $\approx 1.3333\hphantom{{}^{*}}$ FR94 $\frac{10}{7}$ $\approx 1.4286\hphantom{{}^{*}}$ FR94 x $D_{5}$ $\frac{2}{3}$
@@
6 36 $\frac{18}{13}$ $\approx 1.3846$ $\frac{16}{11}$ $\approx 1.4545$ x $E_{6}$ $\frac{2}{3}$
28 $\frac{28}{25}$ $=1.12$ $\frac{133}{109}$ $\approx 1.2202$ ETF LS66 $\frac{39}{24}$
@@
63 $\frac{27}{19}$ $\approx 1.4211$ $\frac{2961}{1991}$ $\approx 1.4872$ x $E_{7}$ $\frac{7}{6}$
@@
8 120 $\frac{45}{31}$ $\approx 1.4516\hphantom{{}^{*}}$ FR94 $\frac{165}{109}$ $\approx 1.5138^{*}$ x $E_{8}$ $\frac{13}{6}$
[…]

### v3

# Better bounds on finite-order Grothendieck constants

@@ Abstract
Grothendieck constants $K_{G}\left({d}{}\right)$ bound the advantage of $d$-dimensional strategies over $1$-dimensional ones in a specific optimisation task. They have applications ranging from approximation algorithms to quantum nonlocality. However, apart from $d=2$, their values are unknown. Here, we exploit a recent Frank-Wolfe approach to provide good candidates for lower bounding some of these constants. The complete proof relies on solving difficult binary quadratic optimisation problems. For $d\in\{3,4,5\}$, we construct specific rectangular instances that we can solve to certify better bounds than those previously known; by monotonicity, our lower bounds improve on the state of the art for $d\leqslant 9$. For $d\in\{4,7,8\}$, we exploit elegant structures to build highly symmetric instances achieving even greater bounds; however, we can only solve them heuristically. We also recall the standard relation with violations of Bell inequalities and elaborate on it to interpret generalised Grothendieck constants $K_{G}\left({d}\mapsto{2}\right)$ as the advantage of complex $d$-dimensional quantum mechanics over real qubit quantum mechanics. Motivated by this connection, we also improve the bounds on $K_{G}\left({d}\mapsto{2}\right)$.
@@ I Introduction
Quantum information is another field where this result is influential: following early observations by Tsirelson Tsi87 , an explicit connection has been established with the noise robustness in Bell experiments Bel64 ; AGT06 . These experiments aim at exhibiting a fascinating property of quantum mechanics in correlation scenarios: nonlocality BCP+14 . The link with Grothendieck’s theorem has led to a surge of interest in the value of $K_{G}\left({3}{}\right)$, the Grothendieck constant of order three. Many works have thus demonstrated increasingly precise lower bounds CHSH69 ; Ver08 ; HLZ+15 ; BNV16 ; DBV17 ; DIB+23 and upper bounds Kri79 ; HQV+17 ; DIB+23 on its value. More recently, a numerical method has also been developed to come up with an even more precise (but not provable) estimate of $K_{G}\left({3}{}\right)$ SM25 .
@@
In this article, we combine the recent projection technique from DIB+23 with the powerful solver developed in DBV17 to obtain better lower bounds on $K_{G}\left({d}{}\right)$ for $3\leqslant d\leqslant 9$. Following FR94 , we also consider symmetric structures in high dimensions emerging from highly symmetric line packings FJM18 to suggest even better bounds on $K_{G}\left({4}{}\right)$, $K_{G}\left({7}{}\right)$, and $K_{G}\left({8}{}\right)$, that we unfortunately cannot prove as they involve optimisation problems that we only solve heuristically. We also consider the generalised Grothendieck constants $K_{G}\left({d}\mapsto{2}\right)$ and interpret them as the advantage of $d$-dimensional quantum mechanics over real qubit quantum mechanics, a fact that was already studied in VP08 ; VP09 . Our bounds on $K_{G}\left({d}{}\right)$ are analytical, while those on $K_{G}\left({d}\mapsto{2}\right)$ strongly rely on numerical methods: we post-process the upper bound on $K_{G}\left({3}\mapsto{2}\right)$ to convert it into an exact result, but the lower bounds remain inaccurate.
@@ II Preliminaries
$\mathrm{SDP}_{d}\left({M}\right)=\max\left\{\sum_{x=1}^{m_{1}}\sum_{y=1}^{m_{2}}M_{xy}\left\langle{a_{x}},{b_{y}}\right\rangle\penalty 10000\ \Big|\penalty 10000\ \forall x\in[m_{1}],\penalty 10000\ a_{x}\in S^{d-1},\penalty 10000\ \forall y\in[m_{2}],\penalty 10000\ b_{y}\in S^{d-1}\right\},$ (1)
where $m]=\{1,\ldots,m\}$ and where $S^{d-1}$ is the $d$-dimensional unit sphere, which reduces to $S^{0}=\{-1,1\}$ for $d=1$. By introducing the set
$\mathrm{SDP}_{d}^{m_{1},m_{2}}=\operatorname{conv}\left\{X\penalty 10000\ \big|\penalty 10000\ X_{xy}=\left\langle{a_{x}},{b_{y}}\right\rangle,\penalty 10000\ \forall x\in[m_{1}],\penalty 10000\ a_{x}\in S^{d-1},\penalty 10000\ \forall y\in[m_{2}],\penalty 10000\ b_{y}\in S^{d-1}\right\},$ (2)
@@
$\mathrm{SDP}_{d}\left({M}\right)=\max\left\{\left\langle{M},{X}\right\rangle\penalty 10000\ \big|\penalty 10000\ X\in\mathrm{SDP}_{d}^{m_{1},m_{2}}\right\}.$ (3)
@@
$K_{G}\left({d}\mapsto{n}\right)=\sup\left\{\frac{\mathrm{SDP}_{d}\left({M}\right)}{\mathrm{SDP}_{n}\left({M}\right)}\penalty 10000\ \Big|\penalty 10000\ M\in\mathbb{R}^{m_{1}\times m_{2}},\quad m_{1},m_{2}\in\mathbb{N}\right\}.$ (5)
The standard Grothendieck constant of order $d$ is obtained when $n=1$, in which case we use the shorthand notation $K_{G}\left({d}\mapsto{1}\right)=K_{G}\left({d}{}\right)$. The infinite-order Grothendieck constant $K_{G}$ initially studied in Gro53 corresponds to the limit $\lim_{d\rightarrow\infty}K_{G}\left({d}{}\right)$, but it is out of the scope of our work, although we briefly mention it in Section V. We refer the reader interested in general aspects and further generalisations of these constants to Bri11 .
@@ III.1 Obtaining facets of the symmetrised correlation polytope
When $n=1$, the set $\mathrm{SDP}_{1}$ is a polytope, called the correlation polytope Pit91 . In BNV16 ; DBV17 ; DIB+23 the method used to solve 1 is to start from a point $P\in\mathrm{SDP}_{d}$ and to derive $M$ as a hyperplane separating $P$ from $\mathrm{SDP}_{1}$. This hyperplane is obtained by solving the projection of $P$ onto the correlation polytope $\mathrm{SDP}_{1}$ via Frank-Wolfe algorithms. We refer to Pok24 for a gentle introduction to FW algorithms, to BCC+25 for a complete review, and to (DIB+23, , Appendix C) for the details of our implementation. Note that some accelerations developed in DIB+23 ; DVP24 and used in this work are now part of the FrankWolfe.jl package BDH+24 .
In particular, the symmetrisation described in DVP24 is crucial, but as it depends on the underlying group and its action of the matrices we consider, exposing the full structure of the symmetrised correlation polytope for each group would be tedious. Instead, we give an brief general definition of this polytope and refer to DVP24 for a detailed example. Given a group $G$ acting on $m_{1}]$ and $m_{2}]$ (by means of signed permutations), the symmetrised correlation polytope $G(\mathrm{SDP}_{1}^{m_{1},m_{2}})$ is the convex hull of the averages of all orbits of the vertices of $\mathrm{SDP}_{1}^{m_{1},m_{2}}$ under the action of $G$. Note that this polytope lives in a subspace of the space invariant under the action of $G$ on matrices of size $m_{1}\times m_{2}$.
In Fig. 1 we illustrate the various geometrical cases that we encounter in this work. Their understanding justifies the procedure that we describe in Algorithm 1 to derive facets of the symmetrised correlation polytope. Importantly, for symmetric instances, the dimension of the ambient space is strictly smaller than $m_{1}m_{2}$. In the following, we consistently denote these facets by $A$, while $M$ will indicate separating hyperplanes without this extra property. Putative facets will also be denoted by $A$, that is, in cases where we rely on heuristic methods to compute $\mathrm{SDP}_{1}\left({A}\right)$.
These heuristic methods play an important role throughout the FW algorithm as they quickly give a good direction to make primal progress. Here we present them in a generalised framework that will turn useful when generalising our algorithm to $K_{G}\left({d}\mapsto{2}\right)$ in Section IV. At a given step of the FW algorithm minimising the squared distance to the set $\mathrm{SDP}_{n}$, finding the best direction with respect to the current gradient $M$ is exactly the problem in Eq. 3. This subroutine is called the Linear Minimisation Oracle (LMO) and, in the course of the algorithm, we usually use an alternating minimisation to obtain heuristic solutions that are enough to make progress (DIB+23, , Appendix B.1). This procedure is known as a seesaw optimisation in the quantum community WW01 .
@@
Figure 1: Illustration of our projection method to separate a matrix $P\in\mathrm{SDP}_{d}$ from the correlation polytope $\mathrm{SDP}_{1}$. (a) In all cases that we consider in this work, the direction $P$ already defines, together with the corresponding value of $\mathrm{SDP}_{1}\left({P}\right)$, a separating hyperplane. (b) When running our FW algorithm to minimise the distance between $P$ and $\mathrm{SDP}_{1}$, the orthogonal projection obtained does not, in general, lie on a facet. The resulting separating hyperplane $M$ is better than $P$ from (a), but still not optimal. Note that this situation can also happen when starting from a point vPvP on the line $[0,P]$ that is not close enough to $\mathrm{SDP}_{1}$ (not depicted here). (c) Even when $v$ is chosen small enough, if our FW algorithm has not fully converged to the orthogonal projection of $vP$, the separating hyperplane $M$ slightly deviates from a facet. (d) For a small enough $v$ and when our FW algorithm has converged, the resulting separating hyperplane is a facet, denoted AA, that is, the tightest separation leading to the best value of $\left\langle{A},{P}\right\rangle/\mathrm{SDP}_{1}\left({A}\right)$.
@@
[…]
@@
5:  $\mathcal{S}\leftarrow\mathrm{BPCG}(vP) \triangleright XX$ is the projection of $vP$ onto $\mathrm{SDP}_{1}^{m_{1},m_{2}}$ obtained with Algorithm 2 from DIB+23
6:  $M\leftarrow vP-X$ $\triangleright$ direction of the separating hyperplane given by the final gradient
7:  $v\leftarrow\left\langle{M},{P}\right\rangle/\mathrm{SDP}_{1}\left({M}\right)$ $\triangleright$ new point on the line $[0,P]$, strictly closer to $\mathrm{SDP}_{1}^{m_{1},m_{2}}$ than the previous one
[…]
9:Output: $A\perp\mathcal{S}$ $\triangleright$ normal vector defining the facet AA of the symmetrisation of $\mathrm{SDP}_{1}^{m_{1},m2}$ BBB+12 ; DVP24
@@ III.2 Heuristic results with highly symmetric line packings
Among the structures that we present in Table 1, the $d$-dimensional configurations reaching the best known kissing number in dimension $d$ are of particular interest. Recall that it corresponds to the greatest number of non-overlapping unit spheres that can be placed such that they each touch (“kiss”) a common unit sphere. In dimension $d=3$, different realisations of the kissing number give rise to different values of $\mathrm{SDP}_{3}\left({A}\right)/\mathrm{SDP}_{1}\left({A}\right)$. As the icosahedron is an optimal line packing, this indicates that giving good bounds on $K_{G}\left({d}{}\right)$ does not, in general, boil down to finding good line packings. Note that the unicity of the configuration in dimension $d=4$ was recently proven LLM24 .
Upon inspection of Table 1 it is apparent that, in the small dimensions that we consider in this work ($d\leqslant 8$), root systems play an important role in obtaining good kissing configurations. These structures are inherently symmetric and play an important role in the theory of Lie groups and Lie algebras. We refer to Bak02 for definitions and elementary properties. The family $D_{d}$ has been studied in FR94 ; there, starting from the Gram matrix $P$ of size d(d1)d(d-1), the exact value of $\mathrm{SDP}_{1}\left({P}\right)$ is computed by invoking symmetry arguments. Moreover, for $d\in\{3,4,5\}$, the optimal diagonal modification of $\lambda=\frac{2}{3}$ is derived for $D_{d}$; in other words, FR94 shows that, among all matrices with shape $P-\lambda\mathds{1}$, the matrix $P-\frac{2}{3}\mathds{1}$ is giving the best lower bound on $K_{G}\left({d}{}\right)$. Here we go a bit further by extending this result to $d\in\{6,7,8\}$, a fact that was suspected in FR94 , but we also observe that the corresponding matrices $A=P-\frac{2}{3}\mathds{1}$ are actually facets of the symmetrised polytope $D_{d}(\mathrm{SDP}_{1}^{d(d-1)})$.
In Table 1, we also give optimal diagonal modifications for other configurations. For the root system $E_{8}$, the value $45/31$ reached without diagonal modification is already provided in (FR94, , page 50), where it is attributed to Reeds and Sloane. It indeed follows from (FR94, , Lemma 2) and from the transitivity of the Weyl group, see the acknowledgements in Section VII. However, these arguments do not directly apply to solve the diagonally modified case.
Note that in all cases involving irrational numbers (except the icosahedron), the facet $A$ cannot be obtained via a diagonal modification of $P$. This is because the off-diagonal elements of $P$ feature rational and irrational numbers. The exception of the icosahedron is due to its extra property of being an equiangular tight frame (ETF), all off-diagonal elements being $\pm 1/\sqrt{5}$, which can then lead to a facet by taking an irrational diagonal modification.
@@
2 3 $\frac{9}{8}$ =1.125=1.125 $\frac{5}{4}$ $1.25\hphantom$$0^{*}$} DVP24 Hexagon $\checkmark$ $A_{2}$ $\frac{2}{3}$
$\frac{9-3\sqrt{5}}{2}$ $\approx 1.1459\hphantom{{}^{*}}$ PV22 $\frac{1+3\sqrt{5}}{6}$ $\approx 1.2847$ Icosahedron $\checkmark$ $H_{3}$ $\frac{15-\sqrt{5}}{15}$
@@
5 20 $\frac{4}{3}$ $\approx 1.3333\hphantom{{}^{*}}$ FR94 $\frac{10}{7}$ $\approx 1.4286\hphantom{{}^{*}}$ FR94 $\checkmark$ $D_{5}$ $\frac{2}{3}$
@@
6 36 $\frac{18}{13}$ $\approx 1.3846$ $\frac{16}{11}$ $\approx 1.4545$ $\checkmark$ $E_{6}$ $\frac{2}{3}$
28 $\frac{28}{25}$ =1.12=1.12 $\frac{133}{109}$ $\approx 1.2202$ ETF LS66 $\frac{39}{24}$
@@
63 $\frac{27}{19}$ $\approx 1.4211$ $\frac{2961}{1991}$ $\approx 1.4872$ $\checkmark$ $E_{7}$ $\frac{7}{6}$
@@
8 120 $\frac{45}{31}$ $\approx 1.4516\hphantom{{}^{*}}$ FR94 $\frac{165}{109}$ $\approx 1.5138^{*}$ $\checkmark$ $E_{8}$ $\frac{13}{6}$
Table 1: Lower bounds on $K_{G}\left({d}{}\right)$ obtained via root systems and other remarkable configurations with symmetry group $G$. Some structures achieve the best known kissing number in their dimension, which we indicate with $\checkmark$. The matrix $A$ is the facet separating the Gram matrix $P$ of the configuration from the symmetrised correlation polytope $G(\mathrm{SDP}_{1})$. When $A=P-\lambda\mathds{1}$, we give the corresponding value of this diagonal modification FR94 in the last column. Shaded cells indicate potential improvements on the state of the art, but in most of these cases, the asterisk indicates that our solution of $\mathrm{SDP}_{1}\left({A}\right)$ is heuristic, hence only putative, see Table 2a. Only the bound on $K_{G}\left({7}{}\right)$ obtained with the $E_{7}$ root system is both exact and good enough to beat the literature; below, by monotonicity of $K_{G}\left({d}{}\right)$, we further improve this bound, see Table 2b. For the 120-cell, we define $\alpha=2566372165103191$, $\beta=1178280120531798$, and $\gamma=10405220765436757$. Note that we are able to exactly obtain $\mathrm{SDP}_{d}\left({P}\right)$ as all these configurations satisfy the platonic property from PV22 . Moreover, we could numerically observe the tightness of our lower bound on $\mathrm{SDP}_{d}\left({A}\right)$ by computing the first level of the Lasserre hierarchy Las01 , see Section IV.2 below.